Name 3. For each of the following production functions below, find the LR total cost as a function of Q: TC(O) LR average total cost as a function of Q: ATCO) .R average total cost vary with Q? (Hint, find the derivative: dATC(Qydo) Does this cost t function exhibit economies of scale, diseconomies of scale, or constant returns to scale? A. Q-4L2 B. Q-2L C. Q-L
Given fixed cost S10 and the total cost given below complete the table assuming fixed price of $8.00. AFC TR Profit AVC ATC TC Total Product MC 20 28 lo 20 -12 lo ID -12 24 36 40 48 58 72 90 110 2o 4 2.5 6,5 9 40 48 I.cc 1.12 8.28 ty 7.259 s0 . 30 10 10
Answer #1

Solution:

Total cost, TC = w*L, where w is wage rate per hour of labor (which is a constant)

Average total cost, ATC = TC(Q)/Q,

partial ATC(Q)/partial Q depicts how average total cost changes with change in Q. If it's negative, this means there exists a negative relation between ATC and Q, that is, with increase in Q, ATC decreases (similarly, positive sign implies a positive relation between the two, and 0 implies no impact of change in Q on ATC that is change in quantity doesn't effect average cost).

Also, economies of scale is achieved when with increase in scale of operation (i.e., production, which Q), cost per unit (which is average cost) decreases. In similar way, diseconomies of scale is achieved when with increase in scale of operation, cost per unit also increases, and constant returns to scale is seen when change in scale of operation doesn't impact cost per unit in any way; then:

A) Q = 4L2

L = (Q/4)1/2 = 0.5*Q0.5

So with, TC = w*L

TC(Q) = w*(0.5*Q0.5) (on substitution)

TC(Q) = 0.5w*Q0.5

ATC(Q) = TC(Q)/Q = 0.5w*Q0.5/Q = 0.5w/Q0.5

partial ATC(Q)/partial Q = -0.5*(0.5w)Q-1.5 = -0.25w/Q1.5

Clearly, with increase in quantity Q, keeping other things unchanged, the average total cost is decreasing (notice the negative sign). This means that this cost function exhibits economies of scale.

B) Q = 2L

L = Q/2 = 0.5*Q

So with, TC = w*L

TC(Q) = w*(0.5*Q) (on substitution)

TC(Q) = 0.5w*Q

ATC(Q) = TC(Q)/Q = 0.5w*Q/Q = 0.5w

partial ATC(Q)/partial Q = 0

Clearly, with increase in quantity Q, keeping other things unchanged, the average total cost is unchanged or constant (since, partial ATC(Q)/partial Q = 0). This means that this cost function exhibits constant returns to scale.

C) Q = L1/3

L = Q3

So with, TC = w*L

TC(Q) = w*(Q3) (on substitution)

TC(Q) = w*Q3

ATC(Q) = TC(Q)/Q = w*Q3/Q = w*Q2

partial ATC(Q)/partial Q = 2*w*Q= 2wQ

Clearly, with increase in quantity Q, keeping other things unchanged, the average total cost is increasing (notice that Q directly increases ATC by a factor of 2w). This means that this cost function exhibits diseconomies of scale.

Q2)

Total product is quantity, Q itself. Given fixed cost (FC), average fixed cost, AFC(Q) = FC/Q. We are given FC = $10 for the exercise. AFC = 10/Q

Total cost, TC(Q) = FC + VC(Q), VC = variable cost, which is a function of Q, that is as Q changes variable cost changes. Thus, at 0 level of output (Q = 0), VC = 0, and TC = FC (=$10 in table above)

Average total cost, ATC = TC/Q

ATC = AFC + AVC , where AVC is average variable cost. Thus, AVC = ATC - AFC

Marginal cost, MC is the additional cost incurred by increasing output by 1 unit. Thus, MC = Tq - Tq-1, q denotes the index of output (in our exercise then, q = {0, 1, 2, 3, ..., 10})

Total revenue, TR = Price*Quantity. With price fixed at $8.00, TR = 8*Q

Finally, profits = total revenue - total cost OR profits = TR - TC

Now, that we have all the required formulas, we can directly fill the table as follows (notice all the formulas are also mentioned in the table alongside the variables to calculate):

AFC= 10/Q ATC TC/Q AVC ATC-AFC MC TC(a) TC(q-1) TR 8* Profits TR-TC 10 Total product, Q TC 10.00 5.00 3.33 2.50 2.00 1.67 1.43 1.25 1.11 1.00 20.00 14.00 11.00 9.00 8.00 8.00 8.29 9.00 10.00 11.00 10.00 9.00 7.67 6.50 6.00 6.33 6.86 7.75 8.89 10.00 12 12 16 24 32 40 48 56 64 72 80 36 48 58 72 90 110 20 30

P.S. your calculations in the table are absolutely correct (except AVC corresponding to Q = 6).